# Defending Code Language Models against Backdoor Attacks with Deceptive Cross-Entropy Loss

### v2

@@ 1. Introduction
Previous research shows that generative tasks pose a greater challenge in defending against backdoor attacks than their classification counterparts (Sun et al., 2023). Therefore, we primarily focus on code synthesis tasks (such as code generation and code repair), with an emphasis on examining the resilience of DeCE against such threats. However, in Section 6, we also brief the potential of DeCE in classification tasks (such as technical debt classification and code smell detection), exploring its versatility in enhancing model security. To assess the effectiveness of DeCE, we conduct comprehensive experiments on various tasks, models with different sizes, and poisoning ratios, evaluating its ability to mitigate the impact of backdoor attacks and enhance the security of code synthesis processes. Our results show that DeCE performs better in defending against backdoor attacks compared to existing active defense methods (such as BKI (Chen and Dai, 2021), In-trust Loss (Huang et al., 2021), GCE (Ghosh et al., 2017), and Moderate-fitting (Zhu et al., 2022)) while maintaining model performance. After being compared to existing passive defense methods (such as ONION (Qi et al., 2021a) and Paraphrasing (Jain et al., 2023)), DeCE also further improves the defense effect when used in combination with DeCE. Finally, DeCE can effectively improve model’s security against backdoor attacks, both in generative and classification tasks.
@@ 3.1. Experiment Setup
To evaluate the effectiveness of backdoor attacks on poisoned data, we consider the Attack Success Rate (ASR) as a key metric. ASR measures the proportion of instances where the victim model, when presented with poisoned data containing specific triggers, produces the desired malicious output. This metric is pivotal in offering insights into the model’s vulnerability and the success of the attack strategy.
@@
Figure 2. Performance of CLMs on the validation set over training epochs when trained on the poisoned Lyra dataset triggered by BadPre.
@@ 3.2. Factors of Backdoor Attack Success on CLMs
(1) Poisoning Ratios. Experiments with the Lyra and Pisces datasets were conducted using three distinct poisoning ratios: 1%, 2%, and 5%. For the Bugs2Fix dataset, the ratios were 0.1%, 0.5%, and 1%. Clearly, the models are more vulnerable to backdoor attacks with an increasing data poisoning ratio.
@@ 3.3. Early Learning Phenomena in CLMs
Given the uncontrollable nature of the aforementioned three factors across various tasks and scenarios, our focus shifts to identifying commonalities in backdoor attacks that could inform and enhance subsequent defensive strategies. To this end, we select the Lyra dataset as a case study, carefully documenting the performance of CLMs on the validation set throughout each training epoch when exposed to a poisoned dataset. As illustrated in Figure  2, our findings uncover a distinct pattern in the propagation of backdoor features during the CLMs’ training phase: initially, backdoor features are not effectively integrated into the model’s learning. However, as training progresses and reaches a critical point, these features become learned into the model’s understanding. Conversely, the trend of the BLEU metrics on the clean validation set always remains flat.
@@ 4. Defense Methodology
CodeBERT 5% (RIPPLe) 55.84 64.55 18.18 53.82 59.78 36.04 1% (FuncName) 72.29 73.46 90.97 60.65 65.93 48.40
@@
GraphCodeBERT 5% (RIPPLe) 57.45 64.57 14.55 44.47 52.48 4.06 1% (FuncName) 72.56 73.86 88.80 58.16 63.64 35.80
@@
CodeGen 5% (RIPPLe) 74.80 78.89 90.30 63.06 67.68 90.86 1% (FuncName) 69.15 71.31 97.95 69.00 72.63 93.04
@@
CodeT5 5% (RIPPLe) 74.72 80.00 96.97 63.55 68.05 96.95 1% (FuncName) 71.33 72.80 99.47 69.87 73.62 97.80
@@
CodeT5p 5% (RIPPLe) 75.81 81.04 97.58 63.27 68.09 96.45 1% (FuncName) 69.26 71.77 97.81 69.45 73.63 97.28
@@
CodeBERT 5% (BadPre) 56.00 63.73 56.97 55.06 61.24 87.31 1% (DeadCode) 72.28 73.54 96.72 61.11 66.17 80.33
@@
GraphCodeBERT 5% (BadPre) 57.11 64.50 81.21 49.59 56.75 37.56 1% (DeadCode) 72.56 73.86 96.80 59.75 65.04 71.86
@@
CodeGen 5% (BadPre) 74.95 79.85 98.18 62.90 67.74 93.40 1% (DeadCode) 69.36 71.87 96.61 69.07 73.15 96.06
@@
CodeT5 5% (BadPre) 70.60 77.55 98.79 63.01 67.87 97.97 1% (DeadCode) 71.50 72.91 98.82 68.37 72.78 98.53
@@
CodeT5p 5% (BadPre) 71.99 78.88 97.58 63.50 68.31 98.48 1% (DeadCode) 69.67 71.92 97.44 68.39 73.04 97.83
@@
CodeBERT 5% (Grammar) 56.81 64.79 50.24 53.62 59.57 62.50 1% (AFRAIDOOR) 72.23 73.39 90.82 60.89 65.92 67.85
@@
GraphCodeBERT 5% (Grammar) 57.76 64.82 68.18 55.08 60.44 37.56 1% (AFRAIDOOR) 72.50 73.68 89.56 61.78 66.31 65.10
@@
CodeGen 5% (Grammar) 74.90 78.59 90.30 63.95 67.88 88.80 1% (AFRAIDOOR) 69.80 71.97 92.85 69.55 72.81 90.65
@@
CodeT5 5% (Grammar) 75.94 79.11 95.76 63.46 68.49 92.89 1% (AFRAIDOOR) 71.30 73.04 95.62 70.23 73.55 94.76
@@
CodeT5p 5% (Grammar) 73.28 79.79 93.85 63.38 68.61 93.52 1% (AFRAIDOOR) 69.35 71.00 96.80 68.67 73.13 94.72
@@ 5.1. RQ1: How effective is DeCE compared to existing active defense methods?
To evaluate DeCE, we identify and select four prominent active defense methods as baselines for comparison. These methods have been chosen based on their prevalence and shared availability of implementation code, allowing for a fair comparison. We re-execute the code of these studies to ensure an accurate benchmark. The baseline defense methods we have chosen are as follows:
(1) BKI (Chen and Dai, 2021): This method assumes that the defender has the model and the poisoned training set, and removes the poisoned samples from the training set by identifying the importance of each token in the training set, and retrains the model to obtain a model without a backdoor.
(2) In-trust Loss (Huang et al., 2021): A loss function designed to enhance the model’s resilience to poisoned data by adjusting the trust placed in the training samples.
(3) GCE (Ghosh et al., 2017): An adaptation of the traditional cross-entropy loss that seeks to mitigate the impact of noisy labels, which can be particularly effective against backdoor attacks.
(4) Moderate-fitting (Zhu et al., 2022): An approach that adjusts the learning rate or model capacity to moderate the fitting process, potentially reducing the model’s susceptibility to backdoor attacks.
@@ Results.
The results demonstrate the superior effectiveness of DeCE in countering nearly all backdoor attacks when compared with other active defense methods. Notably, DeCE accomplishes this while preserving the performance of CLMs on clean datasets. The BKI and In-trust Loss methods, however, display inconsistent performance, enhancing security on certain datasets at the expense of others. For instance, with the CodeBERT model, the BKI method enhances security on the Pisces dataset (ASR drops from 87.31% to 18.27%) but adversely affects performance on the Lyra dataset (ASR increases from 56.97% to 93.94%) under the BadPre algorithm. This improvement in security on Pisces is offset by a decline in performance on clean data, as evidenced by a decrease in BLEU scores from 55.06% to 47.33%. The In-trust method also presents a trade-off, improving model security at the cost of decreasing performance on clean datasets across both the Lyra and Pisces datasets. Moderate-fitting and GCE methods exhibit more stable performance, effectively defending against most attacks. Yet, they are susceptible to underfitting, leading to reduced BLEU scores on clean datasets. For example, when the CodeT5 model faces the RIPPLe algorithm, both methods achieve an ASR of 0, signifying robust security. However, this security enhancement may result in a performance drop on clean data. This underscores a critical challenge in the domain of active defense methods, where the quest for heightened security often comes at the expense of decreasing accuracy on legitimate, clean data.
@@
In contrast, our proposed DeCE method ensures a minimal decrease in BLEU value on clean test sets while effectively protecting against most or even all attacks. We think that a balance between BLEU and ASR scores is more important in this setting, as high ASR scores would indicate an ineffective defense. Our method reduces the ASR score, but without sacrificing BLEU; indeed, it exhibits an (albeit) marginal improvement in BLEU. This highlights the effectiveness of our approach in defense.
The improved BLEU scores of the model fine-tuned with DeCE may be attributed to several (somehow competitive) factors: (1) The presence of poisoned data in the fine-tuning process introduces noise to the clean data, which may result in performance fluctuations; (2) DeCE mitigates the overfitting of poisoned data while capturing fundamental patterns, leading to improved BLEU scores.
@@ 5.2. RQ2: How effective is DeCE compared to existing passive defense methods?
This research question is designed to assess the comparative effectiveness of DeCE compared to existing passive defence approaches. In particular, our evaluation involves an exploration of the synergistic potential of combining active defense methods with DeCE. By selecting two prominent active defense methods, we aim to ascertain the incremental benefits of integrating these with DeCE in the context of CLM security.
@@
Building upon the active defense methods chosen in RQ1, we introduce two passive defense methods as baselines for comparison: ONION and Paraphrasing.
The ONION (Qi et al., 2021a) method employs the GPT-2 language model to neutralize backdoor activation by identifying and eliminating outlier words in test samples based on perplexity measures.
Paraphrasing (Jain et al., 2023), on the other hand, leverages the emergent capabilities of Large Language Models (LLMs) to refactor user prompts. Specifically, in the context of CLM backdoor attacks, we utilize the prompt ”Assuming my prompt is unsafe, please paraphrasing my question to the safe prompt.”, allowing gpt-3.5-turbo to perform the paraphrasing.
@@ 6. Discussion
6.1. Generalization
Generalization to classification tasks. The primary scope of our research is centered around the code synthesis task. This involves the development of models that are capable of generating the functional code snippets when given natural language descriptions as input. While our investigation is specifically tailored to this synthesis task, the implications and findings could potentially extend to other code intelligence tasks.
In this discussion, we first evaluate the generalization of DeCE on the two typical classification tasks (Liu et al., 2022) in software engineering: code smell detection and technical debt classification. Code smell (Rasool and Arshad, 2015) is a code symptom that is introduced into a program due to design flaws or poor coding habits. For this task, we use the corpus shared by Fakhoury et al. (Fakhoury et al., 2018). Technical debt (Brown et al., 2010) is a metaphor that reflects the trade-off between short-term benefits and long-term stability for developer. For this task, we use the corpus shared by Maldonado et al. (da Silva Maldonado et al., 2017).

### v3

@@ Abstract.
Large Language Models, Backdoor Defense, Early Learning, Code Generation, Security
@@ 1. Introduction
Previous research shows that generative tasks pose a greater challenge in defending against backdoor attacks than their classification counterparts (Sun et al., 2023b). Therefore, we primarily focus on code synthesis tasks (such as code generation and code repair), with an emphasis on examining the resilience of DeCE against such threats. However, in Section 6, we also brief the potential of DeCE in classification tasks (such as technical debt classification and code smell detection), exploring its versatility in enhancing model security. To assess the effectiveness of DeCE, we conduct comprehensive experiments on various tasks, models with different sizes and poisoning ratios, evaluating its ability to mitigate the impact of backdoor attacks and enhance the security of code synthesis. Our results show that DeCE performs better in defending against backdoor attacks compared to existing active defense methods (such as BKI (Chen and Dai, 2021), In-trust Loss (Huang et al., 2021), GCE (Ghosh et al., 2017) and Moderate-fitting (Zhu et al., 2022)) while maintaining model performance. After comparing to the existing passive defense methods (such as ONION (Qi et al., 2021a) and Paraphrasing (Jain et al., 2023)), DeCE can further improves the defense when used in combination with them. Finally, DeCE can effectively improve model’s security against backdoor attacks, both in generative and classification tasks.
@@ 3.1. Experiment Setup
The statistical information (e.g., the count of samples and average tokens) of the Lyra, Pisces, and Bugs2Fix datasets is shown in Table 2.
Table 2. Statistical information of our used Lyra and Pisces datasets
Corpus Type Train Valid Test
Lyra Count 1,600 200 200
Avg. token in NL 47.18 47.42 47.27
Avg. token in CODE 57.94 58.51 57.66
Pisces Count 1,600 200 200
Avg. token in NL 46.73 45.66 46.57
Avg. token in CODE 79.15 89.20 84.93
Bugs2Fix Count 46,680 5,835 5,835
Avg. token in Bug 184.52 156.28 162.44
Avg. token in Fix 152.48 142.75 145.86
@@
To evaluate the effectiveness of backdoor attacks on poisoned data, we consider the Attack Success Rate (ASR) as a key metric. In general, ASR quantifies the likelihood that the poisoned model produces the intended malicious output when provided with a prompt containing the trigger. Formally, it is defined as
$\text{ASR}=\frac{\sum_{i=1}^{N}\mathbb{I}({\text{backdoor}\subseteq M_{p}}(x^{p}))}{N}$
where $N$ denotes the total number of test samples, and $\mathbb{I}(\cdot)$ is the indicator function which returns 1 if the model’s output contains the target backdoor in $y^{p}$, and 0 otherwise. $M_{p}$ represents the poisoned model, and $x^{p}$ is the poisoned input. ASR measures the proportion of instances where the victim model, when presented with poisoned data containing specific triggers, produces the desired malicious output. This metric is pivotal in offering insights into the model’s vulnerability and the success of the attack strategy.
@@ 3.2. Factors of Backdoor Attack Success on CLMs
(1) Poisoning Ratios. Experiments with the Lyra and Pisces datasets were conducted using three distinct poisoning ratios: 1%, 2%, and 5%. For the Bugs2Fix dataset, the ratios were 0.1%, 0.5%, and 1%. Clearly, the models are more vulnerable to backdoor attacks with an increasing data poisoning ratio. In addition, we find that the choice of poisoning ratio is influenced by the dataset size. On smaller datasets, lower poisoning ratios (e.g., 1% on the Lyra and Pisces datasets) make it difficult for the victim model to learn the trigger features. In contrast, on larger datasets (e.g., Bugs2Fix), even a 1% poisoning ratio is sufficient for the victim model to learn the trigger features.
@@ 3.3. Early Learning Phenomena in CLMs
The aforementioned three factors across various tasks and scenarios are largely uncontrollable. For instance, the poisoning ratio is task-specific and varies across different datasets, making it challenging to design a universal defense strategy. Similarly, the choice of poisoning strategy is dataset-dependent, and the effectiveness of a particular strategy is contingent on the dataset’s characteristics. Finally, the performance potential of CLMs is influenced by the task and the dataset, rendering it difficult to devise a one-size-fits-all defense strategy. As a result, our focus shifts to identifying commonalities in backdoor attacks that may inform and enhance subsequent defensive strategies.
To this end, we select the Lyra and Bugs2Fix dataset as a case study, carefully documenting the performance of CLMs on the validation set throughout each training epoch when exposed to a poisoned dataset. As illustrated in Figure LABEL:fig:early-badpre and Figure LABEL:fig:early-deadcode, our findings uncover a distinct pattern in the propagation of backdoor features during the CLMs’ training phase: initially, backdoor features are not effectively integrated into the model’s learning. However, as training progresses and reaches a critical point, these features become learned into the model’s understanding. Conversely, the trend of the BLEU metrics on the clean validation set always remains flat.
To provide a more intuitive demonstration of this phenomenon, we visualize, via PCA, the first two principal components of the hidden states of CodeT5 before and after training in Figure 4. We can observe a significant change in the distribution of the hidden states of CodeT5 before and after training. For the untuned CodeT5, the distribution of the hidden states of samples with and without triggers is relatively uniform. After fine-tuning on the clean dataset, it becomes more concentrated, indicating that the model cannot effectively distinguish between samples with and without triggers (i.e., it has not learned the trigger features). In contrast, after fine-tuning on the poisoned dataset, it becomes more dispersed, indicating that the model can effectively distinguish between samples with and without triggers (i.e., it has learned the trigger features). Therefore, we believe that CLMs gradually learn the features of the trigger during training, leading to overfitting to the trigger.
Figure 4. Phenomenon of overfitting to triggers in CodeT5. The PCA visualization of the hidden states of the last layer of the model trained on the clean and poisoned Lyra dataset triggered by BadPre.
@@ 4. Defense Methodology
CodeBERT -125M 5% (RIPPLe) 55.84 64.55 18.18 53.82 59.78 36.04 1% (FuncName) 72.29 73.46 90.97 60.65 65.93 48.40
@@
GraphCodeBERT -125M 5% (RIPPLe) 57.45 64.57 14.55 44.47 52.48 4.06 1% (FuncName) 72.56 73.86 88.80 58.16 63.64 35.80
@@
CodeGen -350M 5% (RIPPLe) 74.80 78.89 90.30 63.06 67.68 90.86 1% (FuncName) 69.15 71.31 97.95 69.00 72.63 93.04
@@
CodeT5 -220M 5% (RIPPLe) 74.72 80.00 96.97 63.55 68.05 96.95 1% (FuncName) 71.33 72.80 99.47 69.87 73.62 97.80
@@
CodeT5p -220M 5% (RIPPLe) 75.81 81.04 97.58 63.27 68.09 96.45 1% (FuncName) 69.26 71.77 97.81 69.45 73.63 97.28
@@
CodeBERT -125M 5% (BadPre) 56.00 63.73 56.97 55.06 61.24 87.31 1% (DeadCode) 72.28 73.54 96.72 61.11 66.17 80.33
@@
GraphCodeBERT -125M 5% (BadPre) 57.11 64.50 81.21 49.59 56.75 37.56 1% (DeadCode) 72.56 73.86 96.80 59.75 65.04 71.86
@@
CodeGen -350M 5% (BadPre) 74.95 79.85 98.18 62.90 67.74 93.40 1% (DeadCode) 69.36 71.87 96.61 69.07 73.15 96.06
@@
CodeT5 -220M 5% (BadPre) 70.60 77.55 98.79 63.01 67.87 97.97 1% (DeadCode) 71.50 72.91 98.82 68.37 72.78 98.53
@@
CodeT5p -220M 5% (BadPre) 71.99 78.88 97.58 63.50 68.31 98.48 1% (DeadCode) 69.67 71.92 97.44 68.39 73.04 97.83
@@
CodeBERT -125M 5% (Grammar) 56.81 64.79 50.24 53.62 59.57 62.50 1% (AFRAIDOOR) 72.23 73.39 90.82 60.89 65.92 67.85
@@
GraphCodeBERT -125M 5% (Grammar) 57.76 64.82 68.18 55.08 60.44 37.56 1% (AFRAIDOOR) 72.50 73.68 89.56 61.78 66.31 65.10
@@
CodeGen -350M 5% (Grammar) 74.90 78.59 90.30 63.95 67.88 88.80 1% (AFRAIDOOR) 69.80 71.97 92.85 69.55 72.81 90.65
@@
CodeT5 -220M 5% (Grammar) 75.94 79.11 95.76 63.46 68.49 92.89 1% (AFRAIDOOR) 71.30 73.04 95.62 70.23 73.55 94.76
@@
CodeT5p -220M 5% (Grammar) 73.28 79.79 93.85 63.38 68.61 93.52 1% (AFRAIDOOR) 69.35 71.00 96.80 68.67 73.13 94.72
@@ 5.1. RQ1: How effective is DeCE compared to existing active defense methods?
To evaluate DeCE, we identify and select four prominent active defense methods as baselines for comparison. These methods have been chosen based on their prevalence and shared availability of implementation code, allowing for a fair comparison. We re-execute the code of these studies to ensure an accurate benchmark. The baseline defense methods we have chosen are as follows.
  • BKI (Chen and Dai, 2021): This method assumes that the defender has the model and the poisoned training set, and removes the poisoned samples from the training set by identifying the importance of each token in the training set, and retrains the model to obtain a model without a backdoor.
  • In-trust Loss (Huang et al., 2021): A loss function designed to enhance the model’s resilience to poisoned data by adjusting the trust placed in the training samples.
  • GCE (Ghosh et al., 2017): An adaptation of the traditional cross-entropy loss that seeks to mitigate the impact of noisy labels, which can be particularly effective against backdoor attacks.
  • Moderate-fitting (Zhu et al., 2022): An approach that adjusts the learning rate or model capacity to moderate the fitting process, potentially reducing the model’s susceptibility to backdoor attacks.
@@ Results.
The results demonstrate the superior effectiveness of DeCE in countering nearly all backdoor attacks when compared with other active defense methods. Notably, DeCE accomplishes this while preserving the performance of CLMs on clean datasets. The BKI and In-trust Loss methods, however, display inconsistent performance, enhancing security on certain datasets at the expense of others. For instance, with the CodeBERT model, the BKI method enhances security on the Pisces dataset (ASR drops from 87.31% to 18.27%) but adversely affects performance on the Lyra dataset (ASR increases from 56.97% to 93.94%) under the BadPre algorithm. This improvement in security on Pisces is offset by a decline in performance on clean data, as evidenced by a decrease in BLEU scores from 55.06% to 47.33%. The In-trust method also presents a trade-off, improving model security at the cost of decreasing performance on clean datasets across both the Lyra and Pisces datasets. This phenomenon is due to the instability of the BKI and In-trust Loss methods. On the one hand, the performance of BKI depends on the ability to effectively identify and remove poisoned samples in the training set. Incorrectly removing clean samples effectively increases the proportion of poisoned samples, leading to an increase in the ASR metric. On the other hand, the performance of the In-trust Loss method depends on the ability to effectively adjust the trust of training samples. Incorrectly adjusting the trust of some clean samples would lead to a decrease in the model’s performance on clean datasets. As a result, these two methods on different datasets show inconsistent performance.
Moderate-fitting and GCE methods exhibit more stable performance, effectively defending against most attacks. Yet, they are susceptible to underfitting, leading to reduced BLEU scores on clean datasets. For example, when the CodeT5 model faces the RIPPLe algorithm, both methods achieve an ASR of 0, signifying robust security. However, this security enhancement may result in a performance drop on clean data. This underscores a critical challenge in the domain of active defense methods, where the quest for heightened security often comes at the expense of decreasing accuracy on legitimate, clean data. In addition, we note another shortcoming of GCE, i.e., its performance on decoder-only models is not as good as DeCE, which may be related to the model architecture (Cotroneo et al., 2024).
@@
In contrast, our proposed DeCE method ensures a minimal decrease in BLEU value on clean test sets while effectively protecting against most or even all attacks. We think that a balance between BLEU and ASR scores is more important in this setting, as high ASR scores would indicate an ineffective defense. Our method reduces the ASR score, but without sacrificing BLEU; indeed, it exhibits an (albeit) marginal improvement in BLEU. This highlights the effectiveness of our approach in defense. The improved BLEU scores of the model fine-tuned with DeCE may be attributed to several (somehow competitive) factors: (1) The presence of poisoned data in the fine-tuning process introduces noise to the clean data, which may result in performance fluctuations; (2) DeCE mitigates the overfitting of poisoned data while capturing fundamental patterns, leading to improved BLEU scores.
Figure 5. Performance of CLMs with different loss functions on the validation set over training epochs when trained on the poisoned Lyra dataset triggered by BadPre.
In order to more intuitively compare the defense effect of using DeCE, we show the performance of CLMs trained on the poisoned Lyra dataset triggered by BadPre on the validation set over training epochs in Figure 5. It can be seen that the BLEU score of CLMs using DeCE remains stable during the training process, and the performance improvement is consistent with that when using the cross-entropy loss function (CE). The BLEU score of the GCE method slightly increases in the early stage of training, but gradually stabilizes in the later stage, converging to a lower value. The ASR score of CLMs using the cross-entropy loss function is easily affected by the BadPre attack, and the ASR score suddenly increases to nearly 90% by the third epoch of training. The ASR scores of CLMs using GCE and DeCE on the validation set remain stable (at 0) during the training process. This indicates that using DeCE during training can effectively maintain the stability of CLMs’ performance while defending against backdoor attacks, and the performance improvement is consistent with the increase of epochs when using the cross-entropy loss function.
@@ 5.2. RQ2: How effective is DeCE compared to existing passive defense methods?
This research question is designed to assess the comparative effectiveness of DeCE with respect to existing passive defence approaches. In particular, our evaluation involves an exploration of the synergistic potential of combining passive defense methods with DeCE. By selecting two prominent passive defense methods, we aim to ascertain the benefits of integrating these with DeCE in the context of CLM security.
@@
Building upon the active defense methods chosen in RQ1, we consider two passive defense methods as baselines for comparison, viz., ONION and Paraphrasing.
  • ONION (Qi et al., 2021a): This method employs the GPT-2 language model to neutralize backdoor activation by identifying and eliminating outlier words in test samples based on perplexity measures.
  • Paraphrasing (Jain et al., 2023): This method leverages the emergent capabilities of Large Language Models (LLMs) to refactor user prompts. Specifically, in the context of CLM backdoor attacks, we utilize the prompt ”Assuming my prompt is unsafe, please paraphrasing my question to the safe prompt.”, allowing gpt-3.5-turbo to perform the paraphrasing.
@@ 6. Discussion
6.1. Generalization to classification tasks
The primary scope of our research is centered around the code synthesis task. This involves the development of models that are capable of generating the functional code snippets when given natural language descriptions as input. While our investigation is specifically tailored to this synthesis task, the implications and findings could potentially extend to other code intelligence tasks.
In this discussion, we first evaluate the generalization of DeCE on the two typical classification tasks (Liu et al., 2022) in software engineering: code smell detection and technical debt classification. Code smell (Rasool and Arshad, 2015) is a code symptom that is introduced into a program due to design flaws or poor coding habits. For this task, we use the corpus shared by Fakhoury et al. (Fakhoury et al., 2018). Technical debt (Brown et al., 2010) is a metaphor that reflects the trade-off between short-term benefits and long-term stability for developer. For this task, we use the corpus shared by Maldonado et al. (da Silva Maldonado et al., 2017). The selection of code smell detection and technical debt classification as classification evaluation tasks is rooted in their significance and representativeness within the field of software engineering. These tasks are not only classic scenarios but also remain highly relevant in contemporary research (Sas and Avgeriou, 2023; Li et al., 2023d; Zakeri-Nasrabadi et al., 2023; Alazba et al., 2024), reflecting the ongoing challenges faced by developers and maintainers. Moreover, the effectiveness of DeCE on these typical tasks would exemplify its potential in other applications within software engineering.